# 3D-COCO: extension of MS-COCO dataset for scene understanding and 3D reconstruction

### v1

@@ 1 Introduction
    COCO 2017 [1]
    80
    80
    164K
    897K
    N/A
    3DObject [14]
    10
    7
    6.7K
    N/C
    N/A
    EPFL Car [15]
    1
    1
    2.3K
    N/C
    N/A
    NYU Depth [16]
    894
    N/C
    1.4K
    35.1K
    N/A
    SUN RGB-D [17]
    $\approx$800
    N/C
    10.3K
    211.2K
    N/A
    KITTI [18]
    2
    2
    15K
    200K
    N/A
    IKEA [19]
    11
    5
    0.8K
    72K
    0.2K
    PASCAL3D+ [20]
    12
    12
    30.9K
    13.8K
    0.1K
    ObjectNet3D [21]
    100
    $\approx$40
    90.1K
    204.6K
    44.1K
    ABO [22]
    63
    $\approx$15
    398.2K
    6.3K
    8K
    ShapeNetCore [2]
    55
    27
    N/A
    N/A
    51.3K
    3D-Future [23]
    34
    $\approx$5
    20.2K
    37.4K
    10.0K
    Google Scans [24]
    $\approx$17
    $\approx$4
    N/A
    N/A
    1K
    CO3D [25]
    50
    50
    1,500K
    18.6K
    18.6K
    Pix3D [26]
    9
    4
    10.1K
    10.1K
    0.4K
    PhotoShape [27]
    1
    1
    N/A
    N/A
    5.8K
    Objaverse [3]
    80
    N/A
    N/A
    $\approx$800K
    Objaverse XL [28]
    80
    N/A
    N/A
    $\approx$10, [^0] K
    3D-COCO (Ours)
    80
    80
    164K
    897K
    28K

### v2

@@ 1 Introduction
    COCO 2017 [1] 80 80 164K 897K N/A
    3DObject [14] 10 7 6.7K N/C N/A
    EPFL Car [15] 1 1 2.3K N/C N/A
    NYU Depth [16] 894 N/C 1.4K 35.1K N/A
    SUN RGB-D [17] $\approx$800 N/C 10.3K 211.2K N/A
    KITTI [18] 2 2 15K 200K N/A
    IKEA [19] 11 5 0.8K 72K 0.2K
    PASCAL3D+ [20] 12 12 30.9K 13.8K 0.1K
    ObjectNet3D [21] 100 $\approx$40 90.1K 204.6K 44.1K
    ABO [22] 63 $\approx$15 398.2K 6.3K 8K
    ShapeNetCore [2] 55 27 N/A N/A 51.3K
    3D-Future [23] 34 $\approx$5 20.2K 37.4K 10.0K
    Google Scans [24] $\approx$17 $\approx$4 N/A N/A 1K
    CO3D [25] 50 50 1,500K 18.6K 18.6K
    Pix3D [26] 9 4 10.1K 10.1K 0.4K
    PhotoShape [27] 1 1 N/A N/A 5.8K
    Objaverse [3] * 80 N/A N/A $\approx$800K
    Objaverse XL [28] *** 80 N/A N/A $\approx$10, [^0] K
    3D-COCO (Ours) 80 80 164K 897K 28K